# Semi-automatic Calculations of Multi-loop Feynman Amplitudes with AmpRed

### v1

@@ Abstract
We present a Mathematica package AmpRed for the semi-automatic calculations of multi-loop Feynman amplitudes with high efficiency and precision. AmpRed implements the methods of integration by parts and differential equations in the Feynman-parameter representation. It allows for the calculations of general parametric integrals (which may not have momentum-space correspondences). Various user-friendly tools for multi-loop calculations, such as those to construct and solve differential equations for Feynman integrals, are provided. It can also deal with tensor algebras in non-relativistic field theories. Interfaces to some packages, like QGRAF and FOMR, are also provided.
@@ 1 Introduction
The first and the most successful technique for integral reduction is the integration-by-parts (IBP) method Tkachov:1981wb ; Chetyrkin:1981qh . Feynman integrals of the same family (integrals with the same set of propagators) are not linearly independent. Instead, they satisfy the IBP identities. All the integrals of the same family can be reduced to a smaller set of integrals (called master integrals) by solving IBP identities according to a prescribed ordering of integrals, a method known as the Laporta algorithm Laporta:2000dsw . It can be shown that the number of master integrals is finite Baikov:2005nv ; Lee:2008tj ; Smirnov:2010hn ; Lee:2013hzt . The Laporta algorithm was implemented in many public packages, including AIR Anastasiou:2004vj , Reduze Studerus:2009ye ; vonManteuffel:2012np , FIRE Smirnov:2008iw ; Smirnov:2014hma ; Smirnov:2019qkx , and Kira Maierhofer:2017gsa ; Klappert:2020nbg . Relying on the Gauss elimination for a large sparse linear-equation system, which leads to large intermediate expressions (referred to as intermediate expression swell), the Laporta algorithm becomes extremely expensive on both time and memory for the reduction of complicated multi-loop integrals, especially those with several scales. The performance can be significantly improved by combining with the finite-field method Kauers:2008zz ; Kant:2013vta ; vonManteuffel:2014ixa ; Peraro:2016wsq ; Klappert:2019emp ; Peraro:2019svx . The efficiency can further be enhanced by refining the IBP systems with techniques like the methods of syzygy equations Gluza:2010ws ; Schabinger:2011dz ; Bohm:2017qme ; Boehm:2020zig ; Wu:2023upw and parametric annihilators Lee:2014tja ; Bitoun:2017nre and the method of block-triangular form Liu:2018dmc ; Guan:2019bcx ; Guan:2024byi . Alternatively to the Laporta-like algorithm, the reduction can be achieved by applying symbolic rules obtained through either the Gröbner-basis technique Baikov:1996iu ; Tarasov:1998nx ; Gerdt:2004kt ; Smirnov:2005ky ; Smirnov:2006tz or solving symbolic IBP identities Lee:2012cn ; Lee:2013mka . Some new developments on integral reduction include the method of intersection theory Mizera:2017rqa ; Mastrolia:2018uzb ; Frellesvig:2019uqt , the method of generating function Feng:2022uqp ; Feng:2022hyg ; Guan:2023avw , etc. .
@@ 2.1 Parametrization
We consider Feynman integrals with the following structure
@@
Here $d=-2\lambda_{0}$, is the spacetime dimension. Conventionally, we define $d=d_{0}-2\epsilon$ with $d_{0}$ the integer dimension. $D_{i}=\sum_{j,k}A_{i,jk}l_{j}\cdot l_{k}+2B_{i,jk}l_{j}\cdot q_{k}+C_{i}$, with $q_{i}$ some external momenta. The $w$ function is defined by Chen:2020wsh   [^2] Note that here the prescription of the infinitesimal imaginary part is different from that in ref. Chen:2020wsh . As a consequence, $w_{0}$ here is different from that in ref. Chen:2020wsh . The IBP identities are insensitive to the imaginary part.
$w_{\lambda}(u)\equiv e^{-\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\mathrm{d}x\,\mathsf{P}\frac{1}{x^{\lambda+1}}e^{iux}~{}.$ (2)
Here we use $\mathsf{P}$ to denote the principal value. We have
$w_{0}(u)=$ $2\pi\left[\theta(u)-\frac{1}{2}\right]~{},$ (3a)
@@
Tensor integrals can be parametrized by using a generator method. That is
@@
where $\widetilde{A}_{ij}\equiv\mathcal{D}_{0}U(A^{-1})_{ij}$ and $\widetilde{B}_{i}^{\mu}\equiv\sum_{j=1}^{L}\widetilde{A}_{ij}B_{j}^{\mu}$.
@@ 2.2 Integral reduction
This method is closer to the standard momentum-space IBP method. In method II, all the indices of the parametric integrals are non-negative, but may be with spacetime dimensions different from that of the original integral in the momentum space.
@@ 2.2.2 Method II
$z^{\prime}_{i}z^{\prime}_{j}$ $i,~{}j=-L-1,~{}-L,\dots,n+1$ (29b)
@@ 2.3 Differential equations
After choosing a specific $\mathcal{E}^{(0)}$, we get the integral $I_{y}$, which can be calculated by using the differential-equation method. The differential-equation system is solved numerically by using the method described in ref. Liu:2022chg . That is, we substitute $\epsilon$ by pure numbers and restore the $\epsilon$ dependence through fitting. The boundary conditions of the differential equations are determined by matching the asymptotic solutions of the differential equations to the asymptotic expansions of the master integrals, which can also be expressed in terms of parametric integrals. Thus the boundary integrals can further be calculated by using the method described in this section. In other words, this method allows us to calculate parametric integrals recursively. The recursive procedure terminates when the $\mathcal{F}$ polynomials of the boundary integrals have at most $n+1$ terms. In this case, the boundary integrals can be evaluated analytically. That is,
@@ 3.2 Feynman integrals
    Similar to the ApartG, it is suggested to divide large expressions into subexpressions while running ToFeynmanInt.
  • AlphaInt[n, {i0,i1,…}] represents a parametric integral of the nth family with indices $\{i_{0},i_{1},\dots\}$.
@@
    "Regulators" is a list of regulators for the indices. By default, "Regulators"->Automatic, indicating that all the symbols appearing in the indices are treated as regulators.
@@
  • AlphaReduce[exp] reduces all the Feynman integrals in exp to master integrals by using the methods described in sec. (2).

### v2

@@ Abstract
We present a Mathematica package AmpRed for the semi-automatic calculations of multi-loop Feynman amplitudes with high efficiency and precision. AmpRed implements the methods of integration by parts and differential equations in the Feynman-parameter representation. It allows for the calculations of general parametric integrals (which may not have momentum-space correspondences). Various user-friendly tools for multi-loop calculations, such as those to construct and solve differential equations for Feynman integrals, are provided. It can also deal with tensor algebras in non-relativistic field theories. Interfaces to some packages, like QGRAF and FORM, are also provided.
@@ 1 Introduction
The first and the most successful technique for multi-loop integral reduction is the integration-by-parts (IBP) method Tkachov:1981wb ; Chetyrkin:1981qh . Feynman integrals of the same family (integrals with the same set of propagators) are not linearly independent. Instead, they satisfy the IBP identities. All the integrals of the same family can be reduced to a smaller set of integrals (called master integrals) by solving IBP identities according to a prescribed ordering of integrals, a method known as the Laporta algorithm Laporta:2000dsw . It can be shown that the number of master integrals is finite Baikov:2005nv ; Lee:2008tj ; Smirnov:2010hn ; Lee:2013hzt . The Laporta algorithm was implemented in many public packages, including AIR Anastasiou:2004vj , Reduze Studerus:2009ye ; vonManteuffel:2012np , FIRE Smirnov:2008iw ; Smirnov:2014hma ; Smirnov:2019qkx , and Kira Maierhofer:2017gsa ; Klappert:2020nbg . Relying on the Gauss elimination for a large sparse linear-equation system, which leads to large intermediate expressions (referred to as intermediate expression swell), the Laporta algorithm becomes extremely expensive on both time and memory for the reduction of complicated multi-loop integrals, especially those with several scales. The performance can be significantly improved by combining with the finite-field method Kauers:2008zz ; Kant:2013vta ; vonManteuffel:2014ixa ; Peraro:2016wsq ; Klappert:2019emp ; Peraro:2019svx . The efficiency can further be enhanced by refining the IBP systems with techniques like the methods of syzygy equations Gluza:2010ws ; Schabinger:2011dz ; Bohm:2017qme ; Boehm:2020zig ; Wu:2023upw and parametric annihilators Lee:2014tja ; Bitoun:2017nre and the method of block-triangular form Liu:2018dmc ; Guan:2019bcx ; Guan:2024byi . Alternatively to the Laporta-like algorithms, the reduction can be achieved by applying symbolic rules obtained through either the Gröbner-basis technique Baikov:1996iu ; Tarasov:1998nx ; Gerdt:2004kt ; Smirnov:2005ky ; Smirnov:2006tz or solving symbolic IBP identities Lee:2012cn ; Lee:2013mka . Some new developments on integral reduction include the method of intersection theory Mizera:2017rqa ; Mastrolia:2018uzb ; Frellesvig:2019uqt , the method of generating function Feng:2022uqp ; Feng:2022hyg ; Guan:2023avw , etc. .
@@ 2.1 Parametrization
We consider scalar Feynman integrals with the following structure
@@
Here $d=-2\lambda_{0}$, is the spacetime dimension. Conventionally, we define $d=d_{0}-2\epsilon$ with $d_{0}$ the integer dimension. $D_{i}=\sum_{j,k}A_{i,jk}l_{j}\cdot l_{k}+2B_{i,jk}l_{j}\cdot q_{k}+C_{i}$, with $q_{i}$ some external momenta. The $w$ function is defined by Chen:2020wsh
$w_{\lambda}(u)\equiv e^{-\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\mathrm{d}x\,\frac{1}{(x-i0^{+})^{\lambda+1}}e^{iux}~{}.$ (2)
We have
$w_{0}(u)=$ $2\pi\theta(u)~{},$ (3a)
@@
Tensor integrals can be parametrized by using a generator method. A tensor integral is obtained by applying a chain of operators $P_{i}^{\mu}$ on a scalar integral. That is,
@@
where $\widetilde{A}_{ij}\equiv\mathcal{D}_{0}U(A^{-1})_{ij}$ and $\widetilde{B}_{i}^{\mu}\equiv\sum_{j=1}^{L}\widetilde{A}_{ij}B_{j}^{\mu}$. The momenta $p_{i}$ are some auxiliary momenta that are absent in the scalar integral $J$. Because both $\tilde{A}$ and $\tilde{B}$ depend on $\mathcal{D}_{0}$, they shift the spacetime dimension of the scalar integral $J$.
@@ 2.2 Integral reduction
This method is closer to the standard momentum-space IBP method. In method II, all the indices of the parametric integrals are non-negative but may be with spacetime dimensions different from that of the original integral in the momentum space.
@@ 2.2.1 Method I
Compared with $\tilde{A}$ and $\tilde{B}$, $\bar{A}$ and $\bar{B}$ are free of $\mathcal{D}_{0}$. Thus they do not shift the spacetime dimension.
@@ 2.2.2 Method II
$z^{\prime}_{i}z^{\prime}_{j},$ $i,~{}j=-L-1,~{}-L,\dots,n+1$ (29b)
@@ 2.3 Differential equations
After choosing a specific $\mathcal{E}^{(0)}$, we get the integral $I_{y}$, which can be calculated by using the differential-equation method. The differential-equation system is solved numerically by using the method described in ref. Liu:2022chg . That is, we substitute $\epsilon$ by pure numbers and restore the $\epsilon$ dependence through fitting. The boundary conditions (chosen at $y=0$) of the differential equations are determined by matching the asymptotic solutions of the differential equations to the asymptotic expansions of the master integrals, which can also be expressed in terms of parametric integrals. Thus the boundary integrals can further be calculated by using the method described in this section. In other words, this method allows us to calculate parametric integrals recursively. The recursive procedure terminates when the $\mathcal{F}$ polynomials of the boundary integrals have at most $n+1$ terms. In this case, the boundary integrals can be evaluated analytically. That is,
@@
2.4 An example
To illustrate how the methods described in this section are carried out in practical calculations, we demonstrate the calculations of the following sunset integral in detail.
$J_{1,12}^{\mu_{1}\mu_{2}}(-\frac{d}{2},\lambda_{1},\lambda_{2},\lambda_{3})=\pi^{-D}\int\mathrm{d}^{d}l_{1}\mathrm{d}^{d}l_{2}\frac{l_{1}^{\mu_{1}}l_{2}^{\mu_{2}}}{\left(l_{1}^{2}\right)^{\lambda_{1}+1}\left(l_{2}^{2}\right)^{\lambda_{2}+1}\left[\left(l_{1}+l_{2}+q\right)^{2}-m^{2}\right]^{\lambda_{3}+1}}~{},$
where $q^{2}=s$. Here the first subscript $1$ in $J_{1,12}$ is just an arbitrary number to distinguish this integral family from others, and the subscripts $12$ correspond to the subscripts of the loop momenta in the numerator.
In the code, this integral can be input by (see sec. 3.2)
In[1]:= int=TI[{l[1], l[2]}, {l[1], 0}, {l[2], 0}, {l[1] + l[2] + p, m},
FV[l[1], a]*FV[l[2], b]];
Here we have replaced the indices $\mu_{1}$ and $\mu_{2}$ by $a$ and $b$ for simplicity.
The $\mathcal{F}$ polynomial of this integral family reads
$\mathcal{F}_{1}~{}=~{}m^{2}\left(x_{1}x_{3}^{2}+x_{2}x_{3}^{2}+x_{1}x_{2}x_{3}\right)-sx_{1}x_{2}x_{3}+\left(x_{1}x_{2}+x_{3}x_{2}+x_{1}x_{3}\right)x_{4}~{}.$
And the matrices $A$, $B$, $\tilde{A}$, and $\tilde{B}$ are
$A_{1}=$ $\begin{pmatrix}\hat{x}_{1}+\hat{x}_{3}&\hat{x}_{3}\\ \hat{x}_{3}&\hat{x}_{2}+\hat{x}_{3}\end{pmatrix}~{},$
$B_{1}^{\mu}=$ $(\hat{x}_{3},\hat{x}_{3})~{}q^{\mu}~{},$
$\tilde{A}_{1}=$ $\mathcal{D}_{0}\begin{pmatrix}\hat{x}_{2}+\hat{x}_{3}&-\hat{x}_{3}\\ -\hat{x}_{3}&\hat{x}_{1}+\hat{x}_{3}\end{pmatrix}~{},$
$\tilde{B}_{1}^{\mu}=$ $\mathcal{D}_{0}(\hat{x}_{2}\hat{x}_{3},\hat{x}_{1}\hat{x}_{3})~{}q^{\mu}~{}.$
2.4.1 Reduction with method II
We consider the reduction with method II first. The integral $J_{1,12}^{\mu_{1}\mu_{2}}$ is parametrized by
$J_{1,12}^{\mu_{1}\mu_{2}}=\left[P_{1}^{\mu_{1}}P_{2}^{\mu_{2}}J_{1}\right]_{p=0}=\left(\tilde{B}_{1,1}\tilde{B}_{1,2}q^{\mu_{1}}q^{\mu_{2}}-\frac{1}{2}\tilde{A}_{1,12}g^{\mu_{1}\mu_{2}}\right)J_{1}~{}.$
$\tilde{B}_{1,1}\tilde{B}_{1,2}$ can further be reduced to
$\tilde{B}_{1,1}\tilde{B}_{1,2}\approx\frac{1}{2s}(\hat{z}_{1}+\hat{z}_{2}-\hat{z}_{3})\left[(1+\hat{a}_{0})\mathcal{D}_{0}U_{1}+1-\hat{a}_{0}\right]+\tilde{B}_{1,1}+\tilde{B}_{1,2}+\frac{m^{2}}{2s}-\frac{1}{2}~{}.$
We specific to the reduction of the integral $J_{1,12}^{\mu_{1}\mu_{2}}(-\frac{d}{2},0,0,0)$. Since all the subsectors of this integral family vanish, we get
$J_{1,12}^{\mu_{1}\mu_{2}}(-\frac{d}{2},0,0,0)=$ $q^{\mu_{1}}q^{\mu_{2}}\left[-\frac{d}{2s}J_{1}(-\frac{d}{2}-1,0,0,1)+J_{1}(-\frac{d}{2}-1,1,0,1)+J_{1}(-\frac{d}{2}-1,0,1,1)\right.$
$\left.+\frac{1}{2s}(m^{2}-s)J_{1}(-\frac{d}{2},0,0,0)\right]+\frac{1}{2}g^{\mu_{1}\mu_{2}}J_{1}(-\frac{d}{2}-1,0,0,1)$
In the code, the above steps are implemented by
In[2]:=AlphaParametrize[int, GroebnerBasis->GroebnerBasis]
The resulting parametric integrals are first reduced by applying symbolic rules generated by eq. (29). In practice, we add a degree bound for efficiency while generating Gröbner basis. Integrals that are not reducible by symbolic rules are $J_{1}(\lambda_{0},0,0,0)$, $J_{1}(\lambda_{0},0,0,1)$, and $J_{1}(\lambda_{0},1,0,0)$. The unreduced integrals are further reduced by solving IBP identities generated by eq. (12). The final result is expressed in terms of two master integrals: $J_{1}(-\frac{d}{2},0,0,0)$ and $J_{1}(-\frac{d}{2}-1,0,0,0)$.
In the code, the reduction is implemented by
In[3]:=AlphaReduce[int]
2.4.2 Reduction with method I
To carry out the reduction with method I, we introduce two auxiliary propagators: $q\cdot l_{1}$ and $q\cdot l_{2}$. The $\mathcal{F}$ polynomial of this integral family reads
$\mathcal{F}_{2}=$ $x_{1}x_{2}x_{3}\left(m^{2}-s\right)+m^{2}\left(x_{1}+x_{2}\right)x_{3}^{2}+\left(x_{1}x_{2}+x_{3}x_{2}+x_{1}x_{3}\right)x_{6}$
$~{}+\frac{s}{4}\left(x_{1}x_{4}^{2}+x_{3}x_{4}^{2}+4x_{1}x_{3}x_{4}-2x_{3}x_{5}x_{4}+x_{2}x_{5}^{2}+x_{3}x_{5}^{2}+4x_{2}x_{3}x_{5}\right).$
The corresponding matrices $\bar{A}$ and $\bar{B}$ are
$\bar{A}_{2}=$ $\begin{pmatrix}-\frac{1}{s}\hat{z}_{5}^{2}-\hat{z}_{1}&\frac{1}{2}\left(m^{2}+\hat{z}_{1}+\hat{z}_{2}-\hat{z}_{3}+2\hat{z}_{4}+2\hat{z}_{5}\right)-\frac{1}{s}\hat{z}_{4}\hat{z}_{5}-\frac{1}{2}s\\ \frac{1}{2}\left(m^{2}+\hat{z}_{1}+\hat{z}_{2}-\hat{z}_{3}+2\hat{z}_{4}+2\hat{z}_{5}\right)-\frac{1}{s}\hat{z}_{4}\hat{z}_{5}-\frac{1}{2}s&-\frac{1}{s}\hat{z}_{4}^{2}-\hat{z}_{2}\end{pmatrix}~{},$
$\bar{B}_{2}^{\mu}=$ $\frac{q^{\mu}}{s}(\hat{z}_{5},\hat{z}_{4})~{}.$
By virtue of eqs. (21 and 22), we have
$J_{1,12}^{\mu_{1}\mu_{2}}(-\frac{d}{2},0,0,0)=$ $J_{2,12}^{\mu_{1}\mu_{2}}(-\frac{d}{2},0,0,0,-1,-1)$
$=$ $\left[\bar{B}_{2,1}\bar{B}_{2,2}q^{\mu_{1}}q^{\mu_{2}}-\frac{1}{2}\tilde{A}_{2,12}\left(g^{\mu_{1}\mu_{2}}-\frac{q^{\mu_{1}}q^{\mu_{2}}}{q^{2}}\right)\right]J_{2}(-\frac{d}{2},0,0,0,-1,-1)$
$=$ $\left[\bar{B}_{2,1}\bar{B}_{2,2}q^{\mu_{1}}q^{\mu_{2}}-\frac{1}{1-d}\bar{A}_{2,12}\left(g^{\mu_{1}\mu_{2}}-\frac{q^{\mu_{1}}q^{\mu_{2}}}{q^{2}}\right)\right]J_{2}(-\frac{d}{2},0,0,0,-1,-1)~{}.$
Since $\bar{A}$ and $\bar{B}$ are free of $\mathcal{D}_{0}$, the resulting parametric integrals are of the same spacetime dimension $d$.
In the code, this step is carried out by
In[4]:=AlphaParametrize[int, Method->1]
The obtained parametric integrals are reduced by solving IBP identities generated by eq. (23), which is implemented in the code by
In[5]:=AlphaReduce[int, Method->1]
2.4.3 Calculation of master integrals
The integrals obtained in the previous subsections can be calculated by constructing differential equations with respect to $m$. Nevertheless, to illustrate the method described in sec. 2.3, we introduce an auxiliary scale $y$ by inserting a delta function $\delta(y-\frac{x_{3}}{x_{4}})$. That is (According to the convention in sec. 2.1, a parametric integral $I$ differs from the corresponding integral $J$ by only a constant prefactor.)
$I_{1}(-\frac{d}{2},0,0,0)=$ $\frac{1}{\Gamma(\lambda_{4}+1)}\int\mathrm{d}y\int\mathrm{d}\Pi^{(4)}~{}\delta(y-\frac{x_{3}}{x_{4}})\mathcal{F}_{1}^{-\frac{d}{2}}x_{4}^{\lambda_{4}}$
$=$ $\frac{1}{\Gamma(\lambda_{4}+1)}\int\mathrm{d}y\int\mathrm{d}\Pi^{(3)}~{}\left(\left.\mathcal{F}_{1}\right|_{x_{3}\to yx_{3},x_{4}\to x_{3}}\right)^{-\frac{d}{2}}x_{3}^{\lambda_{4}+1}$
$\equiv$ $(\lambda_{4}+1)\int\mathrm{d}y~{}I_{3}(-\frac{d}{2},0,0)~{}.$
In the code, this step is implemented by (see secs. 3.2 and 3.3)
In[6]:=int2 = AlphaAddScale[AlphaInt[1, {-D/2, 0, 0, 0, 0}], y]
The obtained $y$-dependent integral can be calculated by using the differential-equation method. For the integral family $I_{3}$, there is only one master integral. The differential equation of it reads
$\frac{\mathrm{d}}{\mathrm{d}y}I_{3}(-\frac{d}{2},0,0)=$ $-\frac{1}{2(3d-2)y\left(m^{2}y+1\right)\left(m^{2}y-sy+1\right)}I_{3}(-\frac{d}{2},0,0)$
×$\times\left[\left(9d^{2}-16d+4\right)m^{2}y^{2}\left(m^{2}-s\right)+2(d-2)(3d-2)\right.$
$\left.+y\left(\left(15d^{2}-32d+12\right)m^{2}+\left(4-3d^{2}\right)s\right)\right]~{}.$
In the code, the differential-equation system is obtained by
In[7]:=des = AlphaDES[AlphaInt[3, {-D/2, 0, 0, 0}], y]
The boundary conditions of the differential-equation system can be determined by matching the asymptotic solution of the differential equation to the asymptotic expansion of the master integral, which are
$I_{3,1}=$ $Cy^{2\epsilon-2}~{},$
$I_{3,2}=$ $y^{2\epsilon-2}I_{4}(-\frac{d}{2},0,0)~{}.$
The $\mathcal{F}$ polynomial of the integral family $I_{4}$ is
$\mathcal{F}_{4}~{}=~{}x_{1}x_{3}^{2}+x_{2}x_{3}^{2}+x_{1}x_{2}x_{3}~{},$
which has exactly $3$ terms. Thus it can be calculated by using eq. (40). That is
$I_{4}(-\frac{d}{2},0,0)~{}=~{}\frac{\Gamma(1-\epsilon)^{2}\Gamma(\epsilon)}{\Gamma(4-3\epsilon)}~{}.$
In the code, the asymptotic solution and the asymptotic expansion are obtained by
In[8]:= asy1 = DESAsymptoticSolve[des[[2]], {y, 0, 0}];
asy2 = AlphaSeries[des[[1, 1]], y -> 0]//AlphaIntEvaluate;
Here AlphaSeries carries out the asymptotic expansion and AlphaIntEvaluate evaluates the obtained integral in terms of gamma functions.
After fixing the boundary conditions, the differential-equation system can be solved either analytically or numerically.
@@ 3.2 Feynman integrals
    Similar to ApartG, it is suggested to divide large expressions into subexpressions while running ToFeynmanInt.
  • AlphaInt[n, $\lambda_{0}$,$\lambda_{1}$,…,$\lambda_{n+1}$}] represents a parametric integral of the nth family with indices $\{\lambda_{0},\lambda_{1},\dots,\lambda_{n+1}\}$, where $\lambda_{i}$ are those appeared in the parametric integral in eq. (4). In eq. (4), the index $\lambda_{n+1}$ is not explicitly specified because it is not independent of the other $\lambda_{i}$. In the code, one may input $\lambda_{n+1}$ with an arbitrary number (it can be automatically corrected). AmpRed label each integral family with an integer n. Information related to an integral family n can be obtained with the function AlphaInfo (see below).
@@
    "Regulators" is a list of regulators for the indices. By default, "Regulators"->
    Automatic
    , indicating that all the symbols appearing in the indices are treated as regulators.
@@
  • AlphaReduce[exp] reduces all the Feynman integrals in exp to master integrals by using the methods described in sec. 2.